# Gradient Boosting Reinforcement Learning

### v1

@@ 6 Conclusion
Historically, RL practitioners have relied on tabular, linear, and NN-based function approximators. But, GBT, a widely successful tool in supervised learning, has been absent from this toolbox. We present a method for effectively integrating it into RL and demonstrate domains where it excels compared to NNs. GBRL is a step toward solutions that are more interpretable, well suited for real-world tasks with structured data, or capable of deployment on low-compute devices.
The choice of an RL method depends on the task characteristics: tabular and linear approaches are suitable for small state spaces or simple mappings, while NNs handle complex relationships in unstructured data. GBT thrives in complex, yet structured environments. In such cases, we observe the advantage of GBRL over NNs, reflecting its already known benefits in supervised learning.
A crucial component of GBRL is our efficient adaptation of GBT for AC methods, which allows the simultaneous optimization of distinct objectives. We optimized this approach for large-scale ensembles using GPU acceleration (CUDA). Furthermore, GBRL integrates seamlessly with existing RL libraries, promoting ease of use and adoption.
7 Limitations and Future Directions
In this work, we integrated the highly popular GBT, typically used in supervised learning, into RL. Our results show that GBT is competitive across a range of problems. However, we identified several limitations and compelling areas for further research. First, a significant challenge lies in the continuous generation of trees. As the policy improves through numerous updates, the size of the ensemble increases. This unbounded growth has implications for memory usage, computational efficiency, and the feasibility of online real-time adaptation. The problem is exacerbated by off-policy methods that build many trees per sample. Moreover, the redundancy of trees, especially those from early stages, suggests that the final policy could be represented with a much smaller ensemble. Consequently, developing strategies for tree pruning, ensemble compression, or dynamically managing ensemble size could offer crucial optimizations without compromising performance.
Another key challenge lies in effectively integrating GBT with additional state-of-the-art RL algorithms such as DDPG [23] or SAC [17]. These require differentiable Q-functions to update the policy. Since GBTs are not differentiable, new solutions are needed to incorporate them into these algorithms. One such possible direction can be probabilistic trees, where each node represents the probability of traversing the graph.
@@ 7 Limitations and Future Directions
This appendix provides supplementary materials that support the findings and methodologies discussed in the main text. It is organized into four sections to present the full experiment results, implementation details, hyperparameters used during the experiments, training progression plots, and experimental plots, respectively. These materials offer detailed insights into the research process and outcomes, facilitating a deeper understanding and replication of the study.
Appendix A Implementaion Details and Hyperparameters
@@
Included in this section are implementation details, information regarding compute resources, and tables containing the hyperparameters used in our experiments enabling the reproducibility of our results. Table 1 lists GBRL hyperparameters for all experiments.
A.1 Environments
@@
The Football domain consists of a vectorized 115-dimensional observation space that summarizes the main aspects of the game and 19 discrete actions. We focus on its academy scenarios, which present situational tasks involving scoring a single goal. A standard reward of +1 is granted for scoring, and we employed the "Checkpoints" shaped reward structure. This structure provides additional points as the agent moves closer towards the goal, with a maximum reward of 2 per scenario. The Atari-ram environment consists of a vectorized 128-dimensional observational space representing the 128 byte RAM state and up to 18 discrete actions. We trained agents in both domains for 10M timesteps.
The MiniGrid environment [8] is a 2D grid world with goal-oriented tasks requiring object interaction. The observation space consists of a 7x7 image representing the grid, a mission string, and the agent’s direction. Each tile in the observed image contains a 3D tuple dictating an object’s color, type, and state. All MiniGrid tasks emit a reward of +1 for successful completion and 0 otherwise.
@@
We trained our NN-based agents on a flattened observation space using the built-in one-hot wrapper. For GBRL agents, we generated a 51-dimensional categorical observational space by encoding each unique tile tuple as a categorical string to represent the observed image. Categorical features were added for the agent’s direction (up, left, right, down) and missions. All agents were trained for 1M timesteps, except for PutNear, FourRooms, and Fetch tasks, which were trained for 10M based on the reported values for PPO NN in RL Baselines3 Zoo.
@@
(a) PPO. For continuous action spaces we used log std init = -2 and log std lr = lin_0.0017. We utilized gradient norm clipping for Gym environments. Specifically, 10 for the value gradients and 150 for the policy gradients.
@@ A.2 Compute Resources
Appendix B Detailed Results Tables
This section contains tables presenting the mean and standard deviation of the average episode reward for the final 100 episodes within each experiment. More specifically, Table 2 presents results for Continuous Control & Block2D environments, Tables 3 and 4 present results for the high-dimensional vectorized environments, and Table 5 presents results for the categorical environments.
Table 2: Continuous-Control and Box2D environments: Average episode reward for the final 100 episodes.
@@ Appendix B Detailed Results Tables
Table 3: Football Academy environments: Average episode reward for the final 100 episodes.
@@
Table 4: Atari-ramNoFrameskip-v4 environments: Average episode reward for the final 100 episodes.
@@
Table 5: MiniGrid environments: Average episode reward for the final 100 episodes.
@@
Appendix C Training Plots
This section presents learning curves depicting model performance throughout the training phase. Figures 8, 9, 10 and 11 show the training reward as a function of environment steps of the agents trained in the experiments. The column order is: A2C, AWR, and PPO.
Figure 8: Classic Control and Box2D environments: Training reward as a function of environment steps.
Figure 9: Football Academy environments: Training reward as a function of environment step.
Figure 10: Atari-ramNoFrameskip-v4 environments: Training reward as a function of environment step.
Figure 11: MiniGrid environments: Training reward as a function of environment step.

### v2

@@ 8 Conclusion
Historically, RL practitioners have relied on tabular, linear, and NN-based function approximators. While GBT evolved as a widely successful tool in supervised learning, it has been absent from this toolbox. In this work, we introduced a method for effectively integrating GBT into RL. We demonstrated domains where GBT excels compared to NNs, in addition to analyzing its various inherent characteristics.
We observe that the ideal method depends on the task characteristics. Tabular and linear approaches are suitable for small state spaces or simple mappings, while NNs excel when tackling complex relationships in unstructured data. In comparison, GBT thrives in complex, yet structured environments. In such cases, GBRL’s advantage is highlighted, reflecting GBTs known benefits from supervised learning.
A crucial component of GBRL is our efficient adaptation of GBT for Actor-Critic methods, which allows the simultaneous optimization of distinct objectives. We optimized this approach for large-scale ensembles using GPU acceleration (CUDA). Furthermore, GBRL integrates seamlessly with existing RL libraries, promoting ease of use and adoption.
GBRL is a step toward well-suited solutions for real-world tasks with structured data.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
This appendix provides supplementary materials that support the findings and methodologies discussed in the main text. It is organized into five sections to present the full experiment results, implementation details, hyperparameters used during the experiments, training progression plots, and experimental plots, respectively. These materials offer detailed insights into the research process and outcomes, facilitating a deeper understanding and replication of the study.
Appendix A Multi-Objective Training With GBT
In GBRL, we compute gradients for both the actor and critic objectives per timestep. We then concatenate these gradients: $g_{t}:=[g_{t,\text{actor}},g_{t,\text{critic}}]$ and use them to build a decision tree. At each candidate node split, we evaluate a score function to decide the best split. However, since the actor and critic gradients can differ in magnitude, this score can become biased toward one objective. To address this, we explored two strategies:
  • Gradient normalization per output dimension with L2-based split scoring.
  • Cosine similarity, which emphasizes gradient direction over magnitude.
Both approaches are supported in our codebase and produce good results, while the cosine similarity performed slightly better (Figure 14).
Figure 14: Multi-objective gradient split strategies in PPO-GBRL on CartPole-v1. We compare gradient normalization and cosine similarity as methods for handling actor-critic gradient imbalance when building trees. Both strategies yield good performance, with cosine similarity achieving slightly faster convergence. Curves show average episodic reward over 1M timesteps.
Appendix B Why Do Certain Algorithms Perform Differently Than Others?
Apart from inherent algorithmic differences, the main difference for GBRL is the number of fitted trees per update. This is analogous to the number of update steps in a neural network, since each tree represents a functional gradient step for the actor or critic. To illustrate this effect, we trained agents in the CartPole-v1 environment for 1M timesteps and varied the number of epochs while keeping the batch size and the rollout length fixed, thereby varying the number of updates per rollout. We repeated this experiment for several batch sizes while keeping the rollout length fixed at 2048. As shown in Figure 15, increasing the number of trees added per rollout improves performance, but increases computational cost. Moreover, a larger batch size also increases computational time, due to a larger number of samples used to build each tree. Hence, there is a clear trade-off between computational cost and performance.
This trade-off also explains the difference in performance of various algorithms when combined with GBRL, as their update schemes inherently affect how many trees are fitted per rollout.
  • A2C - performs a single gradient step per rollout using the entire batch. In GBRL, this results in a single constructed tree per rollout. Since GBTs update the value and policy functions directly, such large, coarse updates can destabilize learning.
  • PPO - in contrast, uses minibatches and multiple epochs per rollout to better control the trade-off. This allows GBRL to fit several small trees per rollout, yielding more stable and incremental updates.
  • AWR - is an off-policy algorithm. Therefore, GBRL could potentially build many trees without adding sample complexity. However, we noticed that our best performing models had a huge computational cost to finish training. As a result, we resorted to reducing the number of gradient steps per update at the expense of performance. Future work could investigate methods for optimizing GBRL with a fixed tree budget (such as pruning and distillation), which may enable the applicability of more demanding RL algorithms, such as AWR, using GBRL.
In general, as each tree approximates a gradient step, GBRL benefits from frequent updates to refine its ensemble gradually. This makes it particularly well-suited to algorithms that support minibatching and multiple epochs per rollout.
Figure 15: Effect of number of gradient updates per rollout on PPO-GBRL’s performance in CartPole-v1. We fix the batch size and the number of samples collected per rollout, while varying the number of PPO epochs, thereby changing the number of gradient updates per rollout. This improves sample efficiency, with more epochs leading to faster convergence. Results are shown over 1M timesteps. The number of samples collected per rollout was fixed to 2048 for all experiments.
Appendix C Implementation Details and Hyperparameters
@@
This section provides implementation details, information regarding compute resources, and tables containing the hyperparameters used in our experiments enabling the reproducibility of our results. Table 2 lists GBRL hyperparameters for all experiments.
C.1 Algorithm Details
Our GBRL method is implemented within the Stable Baselines3 framework (Raffin et al., 2021), a widely used library for standardized reinforcement learning algorithms. For comparisons, we use baseline NNs from RL Zoo3 (Raffin, 2020), which provides tuned hyperparameters per environment. For all experiments, we used a standard MLP with two hidden layers, each containing 64 units and tanh activations. The hidden layers are shared between the actor and the critic, and each component has an additional separate linear layer. This architecture is the default in Stable Baselines3 for the environments we tested and is widely used as a baseline in RL literature. We maintained this consistent architecture across all environments to ensure fair comparison.
A2C (Mnih et al., 2016) is a synchronous, on-policy Actor-Critic algorithm designed to improve learning stability. The critic learns a value function, $V(\operatorname{s})$, used to estimate the advantage. This advantage is incorporated into the policy gradient updates, reducing variance and leading to smoother learning. The policy is updated using the following gradient: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(\operatorname{a}|\operatorname{s})A(\operatorname{s},\operatorname{a})]$.
AWR (Peng et al., 2019) is an off-policy Actor-Critic algorithm. Provided a dataset $\mathcal{D}$, AWR updates both the policy and the value through supervised learning. This dataset can be pre-defined and fixed (offline), or continually updated using the agent’s experience (replay buffer). At each training iteration $k$, AWR solves the following two regression problems:
$V_{k}$ $=\operatorname*{arg\,min}_{V}\mathbb{E}_{\operatorname{s},\operatorname{a}\sim\mathcal{D}}[\|G(\operatorname{s},\operatorname{a})-V(\operatorname{s})\|_{2}^{2}]\,,\,$
$\pi_{k+1}$ $=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\operatorname{s},\operatorname{a}\sim\mathcal{D}}[\log\pi(\operatorname{a}|\operatorname{s})\exp(\frac{1}{\beta}A_{k}(\operatorname{s},\operatorname{a}))]\,,$
where $G(\operatorname{s},\operatorname{a})$ represents the monte-carlo estimate or TD($\lambda$) of the expected return (Sutton & Barto, 2018).
Advantage computation: We use Generalized Advantage Estimation (GAE) as proposed by Schulman et al. (2016) to compute the advantage: $A_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}$ , where $\delta_{t}=r_{t}+\gamma V(\operatorname{s}_{t+1})-V(\operatorname{s}_{t})$. The GAE parameters and for each environment are provided in Table 2 of the appendix. Critic: We use on-policy actor-critic algorithms, where the critic predicts the value function (not action-value Q). We use a standard L2 loss between the critic’s prediction and the target return estimate: $L_{\text{critic}}(\theta)=||V_{\theta}(\operatorname{s}_{t})-G_{t}||^{2}$, where $G_{t}$ is either a Monte Carlo return estimate or a bootstrapped n-step return, depending on the specific algorithm (PPO, A2C, or AWR).
C.2 Environments
Classic-Control & Box2D.
We evaluate GBRL against NNs on all available classic-control tasks provided in Gymnasium (Towers et al., 2024) and on LunarLander-v2. This includes three discrete action space environments: Acrobot-v1, CartPole-v1, MountainCar-v0, and two continuous action space environments: Pendulum-v1 and MountainCarContinuous-v0. We trained on both environments for 1M steps (1.5M for LunarLander-v2). For NNs we used the hyperparameters provided in RL Baselines3 Zoo (Raffin, 2020), except for AWR where we followed the hyperparameters described in the paper.
Google Research Football & Atari ram.
@@ C.2 Environments
The Football domain (Kurach et al., 2020) consists of a vectorized 115-dimensional observation space that summarizes the main aspects of the game and 19 discrete actions. We focus on its academy scenarios, which present situational tasks involving scoring a single goal. A standard reward of +1 is granted for scoring, and we employed the ”Checkpoints” shaped reward structure. This structure provides additional points as the agent moves closer towards the goal, with a maximum reward of 2 per scenario. The Atari-ram domain consists of a vectorized 128-dimensional observational space representing the 128-byte RAM state and up to 18 discrete actions. We trained agents in both domains for 10M timesteps.
MiniGrid.
@@
The MiniGrid environment (Chevalier-Boisvert et al., 2023) is a 2D grid world with goal-oriented tasks requiring object interaction. The observation space consists of a 7x7 image representing the grid, a mission string, and the agent’s direction. Each tile in the observed image contains a 3D tuple dictating an object’s color, type, and state. All MiniGrid tasks emit a reward of +1 for successful completion and 0 otherwise. We trained our NN-based agents on a flattened observation space using the built-in one-hot wrapper, as specified by RL Baselines3 Zoo. For GBRL agents, we generated a 51-dimensional categorical observational space by encoding each unique tile tuple as a categorical string to represent the observed image. Categorical features were added for the agent’s direction (up, left, right, down) and missions. All agents were trained for 1M timesteps, except for PutNear, FourRooms, and Fetch tasks, which were trained for 10M based on the reported values for PPO NN in RL Baselines3 Zoo.
Variable Isolation Environment.
The goal of the variable isolation task is to separate a target variable within a linear equation. In the basic task variation, the initial state is a linear equation of the form $ax+b=c$, where $a,b,c$ are digits within the range $[1-9]$ that can be either positive or negative. The action space is multi-discrete, where the agent chooses an action type, the digit to use, and the sign of the digit. The action types are: addition, subtraction, multiplication, and division. The agent receives a reward proportional to the number of steps taken in the episode if it successfully isolates $x$, otherwise the reward is 00.
In the two-variable linear equation variation, the initial state is a linear equation of the form $ax+by+c=d$, and the goal is to isolate $y$ on the left side. The actions are equivalent to the basic task, with the addition or subtraction of a digit multiplied by $x$. Finally, in the balanced two-variable linear equation variation, the initial state is a linear equation of the form $ax+b=cy+d$, and the goal is to isolate $y$ on the left side. The actions are equivalent to the two-variable linear equation variation with the addition of the operations involving $y$. In all the environment variations, the agent observes the coefficients as inputs. For the two variable variations, we add a small reward bonus if the agent successfully moves a variable or a constant to the correct side. We trained GBRL for 15M steps and NNs for 30M steps.
@@
Variable Isolation 128 0.62 0.08 0.95 0.99 20 256 16 0.18 0.006
(a) PPO. For continuous action spaces, we used log std init = -2 and log std lr = lin_0.0017. We utilized gradient norm clipping for Gym environments. Specifically, 10 for the value gradients and 150 for the policy gradients.
@@ C.3 Compute Resources
Appendix D Detailed Result Tables
This section contains tables presenting the mean and standard deviation of the average episode reward for the final 100 episodes within each experiment. More specifically, Table 3 presents results for Continuous Control & Block2D environments, Tables 4 and 5 present results for the high-dimensional vectorized environments, and Table 6 presents results for the categorical environments.
Table 3: Continuous-Control and Box2D environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
@@ Appendix D Detailed Result Tables
Table 4: Football Academy environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
@@
Table 5: Atari-ramNoFrameskip-v4 environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
@@
Table 6: MiniGrid environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
@@
Appendix E Learning Curves
This section presents learning curves depicting model performance throughout the training phase. Figures 16, 17, 18 and 19 show the training reward as a function of environment steps of the agents trained in the experiments. The column order is: A2C, AWR, and PPO.
Figure 16: Classic Control and Box2D environments: Training reward as a function of environment steps. Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
Figure 17: Football Academy environments: Training reward as a function of environment step. Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
Figure 18: Atari-ramNoFrameskip-v4 environments: Training reward as a function of environment step. Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
Figure 19: MiniGrid environments: Training reward as a function of environment step. Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.